Please tell me in as much detail as possible how to solve this equation:
m x S.H. x T
25 x 4.18 (T-298) = 25 x 0.449 (398-T)
Answer #1

25 x4.18[T- 298] = 25 x 0.449[398-T]

divide both side by 25

then 25 on both side will get cancelled
4.18[T-298] = 0.449[398-T]
brackets can be removed by multiplication
4.18.T - 4.18*298 = 0.449*398 - 0.449.T
4.18.T- 1245.64 =178.702-0.449.T
add both side by 0.449T

4.18T + 0.449T - 1245.64 =178.702 -0.449T+0.449T
4.62T = 178.702 + 1245.64 [on adding both side with 1245.64]
4.62T = 1424.342
T= 1424.342/ 4.62 = 308.299 K
*********************************
hope it is clear .Thank you